The period of oscillation for a mass on a spring system is 0.22 s. If m=3.5 kg, what is the spring

constant of the spring?


1
Expert's answer
2021-01-14T10:38:12-0500

We can find the spring constant of the spring from the formula:


"T=2\pi\sqrt{\dfrac{m}{k}},""T^2=4\pi^2\dfrac{m}{k},""k=\dfrac{4\pi^2m}{T^2},""k=\dfrac{4\pi^2\cdot 3.5\ kg}{(0.22\ s)^2}=2855\ \dfrac{N}{m}."

Answer:

"k=2855\ \dfrac{N}{m}."